aasap please

For an invoice amount over € 5.000, - is regarding the Terms of payment agreed: - payable within 10 days with 2% discount - payable within 30 days without deduction a) What is the duration of the free delivery credit? (1 point) b) Which financing profit will be created for cash discount utilization Using a bank loan with an interest rate of 10%? (1 point) c) What is the interest rate of the delivery credit per year?
Answer #1

a) 10 days

b) Cash discount = 5000*2%=100

Int on bank loan = 5000*10%*30/365=41

Financing profit will be credited to cash discount utilization = 41

c) Int rate of delivery credit per year = 10%per year (if 10% given in the question is assumed as per year rate)

Note: It is considered that a year has 365 days.